Two small magnets each of magnetic moment 10A-m^2 are placed in end on position 0.1m apart from their centres. The force acting between them will me?
  1. 0.6N
  2. 0.06N
  3. 0.06*10^7 N
  4. 0.6*10^7 N.
 
[AP ch8 pg 76 Q59]

Grade:12th pass

1 Answers

Unni
5 years ago
Formula for force BT parallel placed magnet f=Mo/4π(6m1m2/r4)
No=4π×10^-7 therefore f=10^-7(6×10×10/(0.1^4)=0.6N